An iron rail is 20m long. How much will it expand when heated from 100c to 500c (linear expansivity of Iron = 1.2x10-5 0c-1)
1
Expert's answer
2020-06-03T12:11:23-0400

Given, "L_0=20m" ,"T_i=100C,T_f=500C",thus "\Delta T=T_f-T_i=400K".

"\Delta L=L-L_0,\& \alpha_L=1.2 \times 10^{-5}K^{-1}" ,

We know that

"L=L_0(1+ \alpha \Delta T)"

Thus,

"L=20(1+1.2\times 10^{-5}\times 400)\\
\implies L =20.096m"

Therefore, expansion of length is "\Delta L=0.096m"